OpenStudy (anonymous):

calc help, arctan(x/2) - ((1)/(2(x^(2)+4)) if possible help by steps.

OpenStudy (klimenkov):

What to do?

OpenStudy (anonymous):

derivative

OpenStudy (klimenkov):

Ok. \[\arctan \frac x 2-\frac{1}{2(x^2+4)}\]Right?

OpenStudy (anonymous):

correct

OpenStudy (klimenkov):

What is the main problem for you?

OpenStudy (anonymous):

first i seperated the two problems into \[(1/2)/(1+(x ^{2}/2^{2})\]

OpenStudy (anonymous):

um the derivate for the second part and also if the steps to find the derivative for arc tan was right.

OpenStudy (klimenkov):

Can you tell me the derivative of \(\frac1{2(x^2+4)}\)?

OpenStudy (anonymous):

i would use the quotient rule for that correct?

OpenStudy (anonymous):

i got -2/3x ( x^(2)+4)^(-2)

OpenStudy (anonymous):

-x not -2/3x

OpenStudy (klimenkov):

\(\left(\frac1{2(x^2+4)} \right)'=\frac 12\left((x^2+4)^{-1}\right)'=\frac 12\cdot(-1)\cdot(x^2+4)^{-2}\cdot2x=-\frac{x}{(x^2+4)^2}\)

OpenStudy (anonymous):

correct

OpenStudy (klimenkov):

and for arctangent please type too

OpenStudy (anonymous):

arc tan(x/2) = (1/2)/(1+(x^(2)/4)

OpenStudy (klimenkov):

thats right. so can you copy and paste the whole answer?

OpenStudy (anonymous):

arctan(x/2) = (1/2)/(1+(x^(2)/4) - x( x^(2)+4)^(-2)

OpenStudy (anonymous):

(1/2)/(1+(x^(2)/4) + x( x^(2)+4)^(-2)

OpenStudy (klimenkov):

Yeah! Very good. Cant understand why do you ask help, if you can make it by yourself.

OpenStudy (anonymous):

i just wanna make sure im doing it right. like my numbers can be off a little at times

OpenStudy (klimenkov):

Yes. You are absolutely right. Practice with this and your mistakes will vanish.

OpenStudy (anonymous):

alright thanks, quick question though to simplify the arc tan all i would need to do is multiply the 1 by the 4 and bring that 4 and also multiply it by the 1/2?

OpenStudy (klimenkov):

Hm. My English is not perfect. Can you type this, that it can be easy for me to understand?

OpenStudy (anonymous):

(1/2)/(1+(x^(2)/4) this can be turned into 2/(4+X^(2))?

OpenStudy (klimenkov):

Yes.

OpenStudy (anonymous):

thanks